# On the emergence of numerical instabilities in Next Generation Reservoir Computing

### v1

@@ 5. The structure of $\Psi$ characterizes its conditioning
We examine the matrix $\Psi$ in detail, observing that its structure resembles Vandermonde-like and Hankel-like matrices — objects that frequently arise in interpolation, signal processing, and numerical linear algebra [OK01, Ols01]. We numerically demonstrate that $\Psi$ can become ill-conditioned under various parameter settings. The underlying causes are analyzed separately: the influence of the maximum polynomial degree, the interplay between time lags and delay dimension, and the length of training data.
@@
In all numerical experiments, the time series is generated using the explicit forward Euler method with step size $h=0.01$, discarding the initial transient time of $10000$ time steps. Moreover, the time series is normalized component-wise to be mapped inside the cube $[-1,1]^{3}$ by applying the transformation $\mathbf{x}_{n,i}\mapsto\frac{\mathbf{x}_{n,i}}{\max_{n}\{\mathbf{x}_{n,i}\}}$ for each component $i=1,\dots,d$. Since we are interested in the order of magnitude rather than the exact values of the condition number, we normalize $\Psi$ such that its columns have unit norm: for each column $\mathbf{u}_{i}$ of $\Psi$, the new columns are of the form $\frac{\mathbf{u}_{i}}{\|\mathbf{u}_{i}\|}$. This normalization is called column weighting [GVL96], which realizes a lower bound for the condition number of the original matrix [vdS69, Bec00].
5.1. Condition number grows exponentially with respect to $p$
Figure 3. Exponential growth wrt to maximum degree. Box plot of the condition number $\kappa(\hat{\Psi})$ over 25 distinct initial conditions for increasing maximum degree $p$. The exponential growth is confirmed by the exponential curve (black dashed line) plotted for reference. The numerical integrator is the explicit forward Euler method with step size $h=0.01$, and the number of training data points $N_{\mathrm{train}}=5000$.
The first numerical experiment increases the NGRC model’s maximum degree and evaluates the condition number $\kappa(\hat{\Psi})$ over different initial conditions. Fig. 3 shows that $\kappa(\hat{\Psi})$ increases exponentially with respect to $p$, with a growth rate approximately scaling as $e^{3p}$, as confirmed by the linear fit on the logarithmic scale. This suggests that, for systems requiring NGRC models with high-degree polynomials, the associated matrix $\Psi$ is unavoidably ill-conditioned.
The $\Psi$ structure can be related to a well-known matrix in numerical analysis and polynomial interpolation [Hig02]. Note that $\Psi$ evaluates multivariate polynomials along the time series, yielding a Vandermonde-like structure [KRS19]. For example, for $k=1$, $d=2$ and $p=2$:
$\Psi=\begin{pmatrix}1&x_{0,1}&x_{0,1}^{2}&x_{0,2}&x_{0,2}^{2}&x_{0,1}x_{0,2}\\ 1&x_{1,1}&x_{1,1}^{2}&x_{1,2}&x_{1,2}^{2}&x_{1,1}x_{1,2}\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 1&x_{N_{\mathrm{train}}-1,1}&x_{N_{\mathrm{train}}-1,1}^{2}&x_{N_{\mathrm{train}}-1,2}&x_{N_{\mathrm{train}}-1,2}^{2}&x_{N_{\mathrm{train}}-1,1}x_{N_{\mathrm{train}}-1,2}\end{pmatrix}.$
There is a large body of evidence that Vandermonde matrices tend to be badly ill-conditioned [Pan16]. For the univariate case, the ill-conditioning is a consequence of the monomials being a poor basis for the polynomials on the real line. Monomials become increasingly highly correlated as the degree becomes larger over a finite interval. More specifically, it has been shown that the condition number of Vandermonde matrices for the univariate case has exponential growth with respect to $p$; see [GI87, KRS19]. In particular, the growth rate is at least $\big{(}\frac{2}{p+1}\big{)}^{1/2}(1+\sqrt{2})^{p-1}$ for any choice of real positive or symmetrically distributed points (nodes) [Bec00]; for other lower bounds estimates, see [Hig02]. Although precise results for the multivariate case (as relevant for NGRC) are lacking, our numerical observations are consistent with the exponential growth reported in the univariate case.
5.2. The interplay between time lags and delay dimension
Figure 4. Time lag makes $\Psi$ better conditioned. The left panel corresponds to the histogram of the condition number $\kappa(\hat{\Psi})$ over 100 distinct initial conditions for different combinations of NGRC models. Note that the $k=2$ (in orange), $k=3$ (in green), and $k=4$ (in red) overlap, showing a cluster of the different initial conditions at $10^{16}$. The right panel shows that the condition number monotonically decreases for larger time lag $\tau$. The solid line corresponds to the median, and the shaded areas correspond to $25\%$ and $75\%$ quantiles over 25 different initial conditions. The orange dashed line corresponds to the NGRC model with $k=1$, as shown in the left panel. The inset corresponds to the condition number of the submatrix $\hat{\Psi}_{x}$ for increasing $\tau$, whose decaying behavior is captured by the expression written for positive constant $K$ and exponent $\gamma$; see Example 5.1 for details. The numerical integrator is the explicit forward Euler method with step size $h=0.01$, maximum degree $p=2$, and the number of data points $N_{\mathrm{train}}=5000$.
Differently from the maximum degree $p$, which is a known issue in any polynomial interpolation using a monomial basis [Hig02], the effect of adding time-delayed coordinates and time lag is more specific to time-delayed models, such as the NGRC model. The numerical observation is that $\Psi$ is ill-conditioned for small $\tau$. The left panel of Fig. 4 shows the histogram of $\kappa(\hat{\Psi})$ over $100$ different initial conditions for different $k$. The case $k=1$ (in orange) corresponds to the NGRC model akin to the explicit forward Euler method — which will be the reference value (order $10^{2}$). As the delay dimension $k$ increases, the condition number abruptly shifts to $10^{16}$, indicating that the matrix has become rank deficient. We follow a similar argument of [ZdSC25] to pinpoint the specific reason:
Result. For small values of time lag $\tau$ and step size $h$, $\Psi$ is rank deficient for any delay dimension $k>1$.
Proof. Consider the case $\tau=1$. For $k>1$, there exists a subset of monomials in $\mathcal{P}_{p}^{kd}$ that evaluates the trajectory at consecutive time points, and their corresponding columns in $\Psi$ are linearly dependent on each other. This implies that the minimum singular value is close to zero, and consequently, $\Psi$ is rank deficient. For any univariate monomial in $\mathcal{P}_{p}^{kd}$, the argument is the following: choose two distinct columns of $\Psi$ corresponding to the same monomial evaluated at two consecutive time points. Let us denote such monomial as $\phi:\mathbb{R}^{d}\to\mathbb{R}$. By the explicit forward Euler method in Eq. 10 and $\phi$ being continuous differentiable function, the Mean-Value theorem implies that
@@ 5.2. The interplay between time lags and delay dimension
(16) $\mathbf{u}_{1}=\frac{1}{\sqrt{N_{\mathrm{train}}}}\begin{pmatrix}\phi(\mathbf{x}_{0})\\ \phi(\mathbf{x}_{1})\\ \vdots\\ \phi(\mathbf{x}_{N_{\mathrm{train}}-1})\end{pmatrix}\quad\text{and}\quad\mathbf{u}_{2}=\frac{1}{\sqrt{N_{\mathrm{train}}}}\begin{pmatrix}\phi(\mathbf{x}_{1})\\ \phi(\mathbf{x}_{2})\\ \vdots\\ \phi(\mathbf{x}_{N_{\mathrm{train}}})\end{pmatrix}.$
If any two columns of $\Psi$ are equal, it implies that $\Psi$ is singular, i.e., the minimum singular value $\sigma_{m}(\Psi)=0$. Consequently, by continuity of the singular values in terms of the entries of $\Psi$, $\mathbf{u}_{2}=\mathbf{u}_{1}+\mathcal{O}(h)$ with small $h$ implies that $\sigma_{m}(\Psi)$ is close to zero. A similar argument can be made for $\hat{\Psi}$, and other forward numerical integration schemes applied to the original differential equations. In agreement with the left panel of Fig. 4, the decreasing order of singular values of $\Psi$ has a clear gap, confirming that the matrix is rank deficient - results not shown.
We observe that the submatrices formed by evaluating the same monomial at different time-delays resemble a Hankel-like structure, which has already been analyzed in the context of data-driven dynamical systems in Dynamic mode decomposition [AM17]:
@@
5.2.1. Increasing the time lag improves the conditioning of $\Psi$
Although the exponential growth with respect to the maximum degree is inevitable, the ill-conditioning due to small $\tau$ vanishes for large time lags. For a fixed maximum degree and increasing the time lag makes $\Psi$ better conditioned, see the right panel of Fig. 4. It shows that for increasing the time lag, $\kappa(\hat{\Psi})$ decreases and asymptotically approaches the reference value (orange line), which corresponds to the condition number of the NGRC model with $k=1$ and $p=2$ as used in the left panel of Fig. 4. Increasing the time lag eliminates the spurious linear dependence between columns evaluating consecutive time points, which has also been observed in magnetic pendulum dynamics [ZdSC25].
The monotonic decay with respect to the time lag can be related to the statistical properties of the original dynamics: computing the condition number requires to compute the singular values of $\hat{\Psi}$, which are defined to be the square root of eigenvalues of $\hat{\Psi}^{\top}\hat{\Psi}$. First note that $\Psi^{\top}\Psi$ corresponds to the Euclidean inner product of columns $\mathbf{u}_{i}$ of $\Psi$, as in Eq. 16, where each entry is given by
(17) $\begin{split}(\Psi^{\top}\Psi)_{ij}=\langle\mathbf{u}_{i},\mathbf{u}_{j}\rangle&=\frac{1}{N_{\mathrm{train}}}\sum_{n=0}^{N_{\mathrm{train}}-1}\psi_{i}(\mathbf{X}_{n})\psi_{j}(\mathbf{X}_{n})\\ &=\frac{1}{N_{\mathrm{train}}}\sum_{n=0}^{N_{\mathrm{train}}-1}(\psi_{i}\psi_{j})\circ\mathbf{g}_{k,\tau}(\mathbf{f}^{n}(\mathbf{x}_{0}))\\ &=:\frac{1}{N_{\mathrm{train}}}S_{N_{\mathrm{train}}}((\psi_{i}\psi_{j})\circ\mathbf{g}_{k,\tau})(\mathbf{x}_{0}),\end{split}$
@@ 5.2.1. Increasing the time lag improves the conditioning of $\Psi$
Since we are dealing with a chaotic system, it naturally calls for evoking its statistical properties. Let $\mu$ be an invariant ergodic measure preserved under $\mathbf{f}$. By the Birkhoff Ergodic theorem [VO16]: for any $\phi\in L^{1}(\mu)$:
[…]
@@
For sufficiently chaotic dynamics, which is commonly captured by the decay of correlations over time, consider the case when the original dynamics is exponential mixing: $(\mathbf{f},\mu)$ is exponential mixing, such that the correlation function for any $\phi,\varphi,(\phi\cdot\varphi)\in L^{1}(\mu)$ decays:
[…]
@@
For strong chaos, it is commonly associated with the exponential decay of correlations [Via97], including the Lorenz system 63 [AM16].
@@
The Birkhoff Average theorem is valid for the observable $(\psi_{i}\psi_{j})\circ\mathbf{g}_{k,\tau}$ if the pushforward probability measure $(\mathbf{g}_{k,\tau})_{*}\mu$ is an ergodic measure. This depends on the properties of the embedding map $\mathbf{g}_{k,\tau}$, which in turn depends on the delay dimension and time lag. Alternatively, since $\mathbf{g}_{k,\tau}$ evaluates monomial basis at different time, let us denote simply the monomial basis functions on the variables in $\mathbb{R}^{d}$ as $\phi_{i}:M\to\mathbb{R}$, i.e., $\phi_{i}\in\mathcal{P}_{p}^{d}$. Hence, the Birkhoff average in Eq. 17 has a particular form
@@
(20) $\begin{split}\frac{1}{N_{\mathrm{train}}}S_{N_{\mathrm{train}}}((\psi_{i}\psi_{j})\circ\mathbf{g}_{k,\tau})(\mathbf{x}_{0})=\frac{1}{N_{\mathrm{train}}}\sum_{n=0}^{N_{\mathrm{train}}-1}\phi_{i}\circ\mathbf{f}^{q}(\mathbf{f}^{n}(\mathbf{x}_{0}))\phi_{j}\circ\mathbf{f}^{q^{\prime}}(\mathbf{f}^{n}(\mathbf{x}_{0})),\qquad i,j=1,\dots,m,\\ q,q^{\prime}=0,1,\dots,\tau.\end{split}$
By invariance and ergodicity of $\mu$, it follows that for a large set of initial conditions, the Birkhoff average converges to
[…]
@@
For an arbitrary $\mathbf{f}$ and the monomial basis $\mathcal{P}_{p}^{d}$, the right-hand side corresponds to the cross-correlation function for the observables $\phi_{i}$ and $\phi_{j}$ as in Eq. 19. The exponential decay of cross-correlations between delayed observables implies that the spurious linear dependencies among columns of $\Psi$ diminish as $\tau$ increases, as shown by the right panel in Fig. 4. For the monomial basis, it is a hard problem to be able to prove such a result. In contrast, we can study a toy case to grasp this intuition, see Example 5.1.
Example 5.1 ($x$-component case).
@@
Consider the simplest case: $k=2$, $d=1$ and $p=1$, corresponding to $\mathcal{P}_{1}^{2}=\{1,x_{n},x_{n+\tau}\}$. We denote the submatrix formed only with these terms as $\hat{\Psi}_{x}$. Also, assume that the time series $\{x_{n}\}_{n\geq 0}$ has mean zero, which is approximately satisfied by the $x-$component of the Lorenz system due to the system’s symmetry. For this case, $\hat{\Psi}^{\top}\hat{\Psi}$ has the form
@@
(21) $\begin{split}\kappa(\hat{\Psi})&=\sqrt{\frac{1+\langle\mathbf{v}_{2},\mathbf{v}_{3}\rangle}{1-\langle\mathbf{v}_{2},\mathbf{v}_{3}\rangle}}\\ &\leq\sqrt{\frac{1+Ke^{-\gamma\tau}}{1-Ke^{-\gamma\tau}}}=:h_{K,\gamma}(\tau),\end{split}$
@@
for a positive constant $K$ and decay exponent $\gamma$, after applying the exponential decay of correlation bounds in Eq. 19. This formula $h_{K,\gamma}(\tau)$ shows that: while the function can be unbounded for small $\tau$, the function decays to one for large $\tau$. The inset of the right panel in Fig. 4 shows $\kappa(\hat{\Psi}_{x})$ (solid line) for increasing time lag, which decays monotonically to one. The dashed line corresponds to the fitted curve (with $K=1.01$ and $\gamma=0.01$) by the function $h_{K,\gamma}(\tau)$ in Eq. 21 to the numerical value. Interestingly, for $K=1$, the function can be written as $h_{1,\gamma}(\tau)=\coth^{\frac{1}{2}}{\big{(}\frac{\gamma\tau}{2}\big{)}}$.
We observe that the decay of $\kappa(\hat{\Psi})$ is slower than $\kappa(\hat{\Psi}_{x})$, which is related to the presence of the additional monomials. For instance, other observables as the $z-$component have slower correlation decay. A more detailed analysis of the full matrix will be left for future work.
@@
Figure 5. Condition number with respect to the length of training data. Condition number of $\hat{\Psi}$ for increasing $N_{\mathrm{train}}$ for different time steps $h$. The left panel shows the results for the NGRC model with $k=1$ ($m=10$), and the right panel shows the model with $k=2$, $\tau=50$ ($m=28$). In both cases, the NGRC model has a maximum degree $p=2$. All curves for $h>0.1$ overlap for all $N_{\mathrm{train}}$ tested. The solid line corresponds to the median, and the shaded areas correspond to $25\%$ and $75\%$ quantiles over 10 different initial conditions.
5.3. Dependence on the length of training data

### v2

@@ IV The structure of $\Psi$ characterizes its conditioning
By studying time series of chaotic systems, we examine the matrix $\Psi$ in detail, observing that its structure resembles Vandermonde-like and Hankel-like matrices — objects that frequently arise in interpolation, signal processing, and numerical linear algebra [44, 45]. We numerically demonstrate that $\Psi$ can become ill-conditioned under various parameter settings. The underlying causes are analyzed separately: the influence of the maximum polynomial degree, the interplay between time lags and delay dimension, and the length of training data.
@@
In all numerical experiments of this section, the numerical integrator used is the explicit Euler method, discarding the initial transient time of $10000$ time steps. Moreover, since we are interested in the order of magnitude rather than the exact values of the condition number, we normalize $\Psi$ such that its columns have unit norm: for each column $\mathbf{u}_{i}$ of $\Psi$, the new columns are of the form $\frac{\mathbf{u}_{i}}{\|\mathbf{u}_{i}\|}$. This normalization is called column weighting [27], which realizes a lower bound for the condition number of the original matrix [46, 47]. The normalized matrix is denoted $\hat{\Psi}$.
IV.1 Condition number grows exponentially with respect to $p$
@@
[…]
The $\Psi$ structure can be related to a well-known matrix in numerical analysis and polynomial interpolation [29]. Note that $\Psi$ evaluates multivariate polynomials along the time series, yielding a Vandermonde-like structure [48]. For example, for $k=1$, $d=2$ and $p=2$:
${\footnotesize\Psi=\begin{pmatrix}1&x_{0,1}&x_{0,1}^{2}&x_{0,2}&x_{0,2}^{2}&x_{0,1}x_{0,2}\\ 1&x_{1,1}&x_{1,1}^{2}&x_{1,2}&x_{1,2}^{2}&x_{1,1}x_{1,2}\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 1&x_{N_{\mathrm{train}}-1,1}&x_{N_{\mathrm{train}}-1,1}^{2}&x_{N_{\mathrm{train}}-1,2}&x_{N_{\mathrm{train}}-1,2}^{2}&x_{N_{\mathrm{train}}-1,1}x_{N_{\mathrm{train}}-1,2}\end{pmatrix}.}$
There is a large body of evidence that Vandermonde matrices tend to be badly ill-conditioned [49]. For the univariate case, the ill-conditioning is a consequence of the monomials being a poor basis for the polynomials on the real line. Monomials become increasingly highly correlated as the degree becomes larger over a finite interval. More specifically, it has been shown that the condition number of Vandermonde matrices for the univariate case has exponential growth with respect to $p$; see [50, 48]. In particular, the growth rate is at least $\big(\frac{2}{p+1}\big)^{1/2}(1+\sqrt{2})^{p-1}$ for any choice of real positive or symmetrically distributed points (nodes) [47]; for other lower bounds estimates, see [29]. Although precise results for the multivariate case (as relevant for NGRC) are lacking, our numerical observations are consistent with the exponential growth reported in the univariate case.
IV.2 The interplay between time lags and delay dimension
[…]
Result IV.1.
For small values of time lag $\tau$ and step size $h$, $\Psi$ is rank deficient for any delay dimension $k>1$.
Proof.
Consider the case $\tau=1$. For $k>1$, there exists a subset of monomials in $\mathcal{P}_{p}^{kd}$ that evaluates the trajectory at consecutive time points, and their corresponding columns in $\Psi$ are linearly dependent on each other. This implies that the minimum singular value is close to zero, and consequently, $\Psi$ is rank deficient. For any univariate monomial in $\mathcal{P}_{p}^{kd}$, the argument is the following: choose two distinct columns of $\Psi$ corresponding to the same monomial evaluated at two consecutive time points. Let us denote such monomial as $\phi:\mathbb{R}^{d}\to\mathbb{R}$. By the explicit forward Euler method in Eq.˜10 and $\phi$ being continuous differentiable function, the Mean-Value theorem implies that
$\phi(\mathbf{x}_{n+1})=\phi(\mathbf{x}_{n})+h\Big(\int_{0}^{1}\nabla\phi(\mathbf{x}_{n}+s\mathbf{F}(\mathbf{x}_{n}))ds\Big)\cdot\mathbf{F}(\mathbf{x}_{n}),$ (16)
@@ IV.2 The interplay between time lags and delay dimension
$\mathbf{u}_{1}=\frac{1}{\sqrt{N_{\mathrm{train}}}}\begin{pmatrix}\phi(\mathbf{x}_{0})\\ \phi(\mathbf{x}_{1})\\ \vdots\\ \phi(\mathbf{x}_{N_{\mathrm{train}}-1})\end{pmatrix}~\text{and}~\mathbf{u}_{2}=\frac{1}{\sqrt{N_{\mathrm{train}}}}\begin{pmatrix}\phi(\mathbf{x}_{1})\\ \phi(\mathbf{x}_{2})\\ \vdots\\ \phi(\mathbf{x}_{N_{\mathrm{train}}})\end{pmatrix}$. (17)
If any two columns of $\Psi$ are equal, it implies that $\Psi$ is singular, i.e., the minimum singular value $\sigma_{m}(\Psi)=0$. Consequently, by continuity of the singular values in terms of the entries of $\Psi$, $\mathbf{u}_{2}=\mathbf{u}_{1}+\mathcal{O}(h)$ with small $h$ implies that $\sigma_{m}(\Psi)$ is close to zero. ∎
A similar argument can be made for $\hat{\Psi}$, and other forward numerical integration schemes applied to the original differential equations. In agreement with the left panel of Fig.˜6, the decreasing order of singular values of $\Psi$ has a clear gap, confirming that the matrix is rank deficient - results not shown.
We observe that the submatrices formed by evaluating the same monomial at different time-delays resemble a Hankel-like structure, which has already been analyzed in the context of data-driven dynamical systems in Dynamic mode decomposition [51]:
@@
IV.2.1 Increasing the time lag improves the conditioning of $\Psi$
[…]
The monotonic decay with respect to the time lag can be related to the statistical properties of the original dynamics: computing the condition number requires computing the singular values of $\hat{\Psi}$, which are defined to be the square root of eigenvalues of $\hat{\Psi}^{\top}\hat{\Psi}$. First note that $\Psi^{\top}\Psi$ corresponds to the Euclidean inner product of columns $\mathbf{u}_{i}$ of $\Psi$, as in Eq.˜17, where each entry is given by
$\begin{split}(\Psi^{\top}\Psi)_{ij}=\langle\mathbf{u}_{i},\mathbf{u}_{j}\rangle&=\frac{1}{N_{\mathrm{train}}}\sum_{n=0}^{N_{\mathrm{train}}-1}\psi_{i}(\mathbf{X}_{n})\psi_{j}(\mathbf{X}_{n})\\ &=\frac{1}{N_{\mathrm{train}}}\sum_{n=0}^{N_{\mathrm{train}}-1}(\psi_{i}\psi_{j})\circ\mathbf{g}_{k,\tau}(\mathbf{f}^{n}(\mathbf{x}_{0}))\\ &=:\frac{1}{N_{\mathrm{train}}}S_{N_{\mathrm{train}}}((\psi_{i}\psi_{j})\circ\mathbf{g}_{k,\tau})(\mathbf{x}_{0}),\end{split}$ (18)
@@ IV.2.1 Increasing the time lag improves the conditioning of $\Psi$
Since we are dealing with a chaotic system, it naturally calls for evoking its statistical properties. Let $\mu$ be an invariant ergodic measure preserved under $\mathbf{f}$. By the Birkhoff Ergodic theorem [52]: for any $\phi\in L^{1}(\mu)$:
$\lim_{N\to\infty}\sum_{n=0}^{N-1}\phi\circ\mathbf{f}^{n}(\mathbf{x}_{0})=\int_{M}\phi d\mu,\quad\mu-\text{a.e.}$ (19)
@@
For sufficiently chaotic dynamics, which is commonly captured by the decay of correlations over time, consider the case when the original dynamics is exponential mixing: $(\mathbf{f},\mu)$ is exponential mixing, such that the correlation function for any $\phi,\varphi,(\phi\cdot\varphi)\in L^{1}(\mu)$ decays exponentially for $\tau\geq 0$:
$\Big|\int\phi(\varphi\circ\mathbf{f}^{\tau})d\mu-\int\phi d\mu\int\varphi d\mu\Big|\leq K(\phi,\varphi)e^{-\gamma\tau}.$ (20)
@@
Such exponential decay of correlations [53] is a statistical property commonly found in chaotic systems, including the Lorenz system 63 [54].
@@
The Birkhoff Average theorem is valid for the observable $(\psi_{i}\psi_{j})\circ\mathbf{g}_{k,\tau}$ if the pushforward probability measure $(\mathbf{g}_{k,\tau})_{*}\mu$ is an ergodic measure. This depends on the properties of the embedding map $\mathbf{g}_{k,\tau}$, which in turn depends on the delay dimension and time lag. Alternatively, since $\mathbf{g}_{k,\tau}$ evaluates monomial basis at different time, let us denote simply the monomial basis functions on the variables in $\mathbb{R}^{d}$ as $\phi_{i}:M\to\mathbb{R}$, i.e., $\phi_{i}\in\mathcal{P}_{p}^{d}$. Hence, the Birkhoff average in Eq.˜18 has a particular form
@@
$\begin{split}\frac{1}{N_{\mathrm{train}}}S_{N_{\mathrm{train}}}&((\psi_{i}\psi_{j})\circ\mathbf{g}_{k,\tau})(\mathbf{x}_{0})\\ &=\frac{1}{N_{\mathrm{train}}}S_{N_{\mathrm{train}}}\big((\phi_{i}\circ\mathbf{f}^{q})(\phi_{j}\circ\mathbf{f}^{q^{\prime}})\big)(\mathbf{x}_{0}),\end{split}$ (21)
$i,j=1,\dots,m$ and $q,q^{\prime}=0,1,\dots,\tau.$ By Eq.˜19, the Birkhoff average in Eq.˜21 converges $\mu$-a.e. to
$\lim_{N_{\mathrm{train}}\to\infty}\frac{1}{N_{\mathrm{train}}}S_{N_{\mathrm{train}}}\Big((\phi_{i}\circ\mathbf{f}^{q})(\phi_{j}\circ\mathbf{f}^{q^{\prime}})\Big){(\mathbf{x}_{0})}$ $=\int_{M}(\phi_{i}\circ\mathbf{f}^{q})(\phi_{j}\circ\mathbf{f}^{q^{\prime}})d\mu$
$=\int_{M}\phi_{i}(\phi_{j}\circ\mathbf{f}^{(q^{\prime}-q)})d\mu,$
@@
where the last equality holds due to the invariance of $\mu$ under $\mathbf{f}$. For an arbitrary $\mathbf{f}$ and the monomial basis $\mathcal{P}_{p}^{d}$, the right-hand side corresponds to the cross-correlation function for the observables $\phi_{i}$ and $\phi_{j}$ as in Eq.˜20. The exponential decay of cross-correlations between delayed observables implies that the spurious linear dependencies among columns of $\Psi$ diminish as $\tau$ increases, as shown by the right panel in Fig.˜6. For the monomial basis, it is a hard problem to prove such a result. In contrast, we can examine a toy case to gain insight into this intuition, as shown in Example˜IV.2.
Example IV.2 ($x$-coordinate case).
@@
Consider the simplest case: $k=2$, $d=1$ and $p=1$, corresponding to $\mathcal{P}_{1}^{2}=\{1,x_{n},x_{n+\tau}\}$. We denote the submatrix formed only with these terms as $\hat{\Psi}_{x}$. Also, assume that the time series $\{x_{n}\}_{n\geq 0}$ has mean zero, which is approximately satisfied by the $x$-coordinate of the Lorenz system due to the system’s symmetry. For this case, $\hat{\Psi}^{\top}\hat{\Psi}$ has the form
@@
$\begin{split}\kappa(\hat{\Psi})&=\sqrt{\frac{1+\langle\mathbf{v}_{2},\mathbf{v}_{3}\rangle}{1-\langle\mathbf{v}_{2},\mathbf{v}_{3}\rangle}}\\ &\leq\sqrt{\frac{1+Ke^{-\gamma\tau}}{1-Ke^{-\gamma\tau}}}=:h_{K,\gamma}(\tau),\end{split}$ (22)
@@
for a positive constant $K$ and decay exponent $\gamma$, after applying the exponential decay of correlation bounds in Eq.˜20. This formula $h_{K,\gamma}(\tau)$ shows that: while the function can be unbounded for small $\tau$, the function decays to one for large $\tau$. The inset of the right panel in Fig.˜6 shows $\kappa(\hat{\Psi}_{x})$ (solid line) for increasing time lag, which decays monotonically to one. The dashed line corresponds to the fitted curve (with $K=1.01$ and $\gamma=0.01$) by the function $h_{K,\gamma}(\tau)$ in Eq.˜22 to the numerical value. Interestingly, for $K=1$, the function can be written as $h_{1,\gamma}(\tau)=\coth^{\frac{1}{2}}{\big(\frac{\gamma\tau}{2}\big)}$.
[…]
@@
Figure 7: Condition number with respect to the length of training data. Condition number of $\hat{\Psi}$ for increasing $N_{\mathrm{train}}$ for different time steps $h$. The left panel shows the results for the NGRC model with $k=1$ ($m=10$), and the right panel shows the model with $k=2$, $\tau=50$ ($m=28$). In both cases, the NGRC model has a maximum degree $p=2$. All curves for $h>0.1$ overlap for all $N_{\mathrm{train}}$ tested. The solid line corresponds to the median, and the shaded areas correspond to $25\%$ and $75\%$ quantiles over 10 different initial conditions.
IV.3 Dependence on the length of training data